# Vid2Fluid: 3D Dynamic Fluid Assets from Single-View Videos with Generative Gaussian Splatting

### v1

@@ 4.5. Different Rendering Options
The output 3D Gaussians dynamic assets can be rendered mainly in two ways: retaining the originally generated spherical harmonic features (like the mountain river example rendered in Fig. 11 bottom) or reconstructing meshes for high-resolution rendering and interaction (the mountain river example in Fig. 1 top). Without assigning uniform appearances to fluid, 3D Gaussian rasterization relies on the unknown spherical harmonic features of inlet particles. Thus, a natural choice is to sample features from particles at the initial position. We note that the generative method often produces Gaussians at low resolution, which may limit the application of harmonic features, depending on the fluid type.
@@
Unlike a closed system, our fluid assets include inlet particles, which complicates the prediction of inflow textures. While the appearance of the fluid typically aligns with uniform flow, it may lack consistency with the reconstructed appearance. To enable interaction between the generated Gaussians and other assets, users might need to re-render the fluid using the same illumination pipeline as the rest of the scene. For compatibility with splatting rendering, we could adapt the existing 3D Gaussian relighting frameworks (Jiang et al., 2024) and implement the Gaussian fluid with transparent light absorption  (Feng et al., 2024), ensuring rendering efficiency with 3DGS. Moreover, with current GPU-accelerated parallelization, the simulation process achieves real-time speeds, enabling both rasterization and simulation of Gaussians on the fly.
4.6. User Editing
@@
For users looking to edit the generated fluid asset, the framework provides easy modification of simulation parameters to achieve different fluid effects, change the fluid’s material, adjust the downstream flow, or introduce solid-liquid interactions as shown in Fig. 12. In this way, fluid behavior within scenic videos can be faithfully reproduced in the digital world.
4.7. Runtime
@@
Figure 9. Runtime (in second) of the example shown in Fig. 1 (b).
5. Conclusions
In this paper, we introduced a novel two-stage pipeline for reconstructing consistent and accurate fluid dynamics from a single-view video, starting with geometry and motion reconstruction and then optimizing simulation parameters. This approach has opened up new possibilities in the field of fluid asset creation. However, our approach also has certain limitations. Firstly, it currently cannot handle long videos with high-torrent fluids due to increased complexity and convergence issues in simulation parameter optimization. Color changes not related to velocity (e.g., in lava due to temperature) may also affect the accuracy of our optical-flow-based velocity estimation. In addition, we note that an optional requirement for our method is that camera parameters are needed for single-view videos with a moving camera. Future work will aim to address these limitations for more accurate reconstructions.

### v2

@@ 3.5 Result Rendering
The output 3D Gaussians dynamic assets can be rendered mainly in two ways: retaining the originally generated spherical harmonic features or reconstructing meshes for high-resolution rendering and interaction. Without assigning uniform appearances to fluid, 3D Gaussian rasterization relies on the unknown spherical harmonic features of inlet particles. Thus, a natural choice is to sample features from particles at the initial position. We note that the generative method often produces Gaussians at low resolution, which may limit the application of harmonic features, depending on the fluid type. Through the experiment, rendering directly by splatting is best suited for viscous fluids or fluids with uniform appearance. The former type of fluid flows more steadily like a deformable soft body, and the latter type is free of appearance inconsistency in fluid motion.
@@
Unlike a closed system, our fluid assets include inlet particles, which complicates the prediction of inflow textures. While the appearance of the fluid typically aligns with uniform flow, it may lack consistency with the reconstructed appearance. To enable interaction between the generated Gaussians and other assets, users might need to re-render the fluid using the same illumination pipeline as the rest of the scene. For compatibility with splatting rendering, we could adapt the existing 3D Gaussian relighting frameworks [JTL24] and implement the Gaussian fluid with transparent light absorption  [FFS25], ensuring rendering efficiency with 3DGS. Moreover, with current GPU-accelerated parallelization, the simulation process achieves real-time speeds, enabling both rasterization and simulation of Gaussians on the fly.
Figure 9: Examples of our output. Left column: frames from the input videos. Middle column: frames from the rendered output. Right column: side views of the output fluid, captured from camera angles indicated by the yellow arrows in the middle images. The last two rows correspond to viscous fluids and they are directly rendered by splatting.
3.6 Editable and interactive features
Unlike 4D reconstruction methods based on canonical and displacement fields, recovering fluid dynamics using a physics-based approach offers enhanced editability and interaction with the generated geometry. An optimized set of physical parameters provides an explicit representation of fluid motion, making it more interpretable for artists and users without specialized expertise. In contrast, directly optimized displacement fields typically possess a high degree of freedom, which complicates the interpretation of their influence on fluid behavior. In our method, any modifications to the optimized parameters or external force interactions are propagated through fluid re-simulation, ensuring all changes remain physically consistent. This capability is particularly advantageous when scene reconstruction is not the sole objective (e.g., when extracting 3D assets from real-world videos and integrating them into digital scenery). Moreover, our approach not only recovers the fluid body but also generates the surrounding and underlying terrain, providing enriched geometric information for diverse post-processing applications.
4 Results
We implement the framework of our method mainly using Python on a workstation PC equipped with a 24-core Intel(R) Xeon(R) Platinum 8255C CPU and an NVIDIA Tesla V100 GPU. For the 3DGS generation part, we port the open-source implementation of the existing methods (we use TriplaneGaussian [ZYG24] by default), and output intermediate states. For re-simulating the flowing fluid, we implemented a GPU-parallelized APIC method with Taichi [HLA19]. The rendering of obtained 3D Gaussians is compatible with a general 3DGS visualizer. We also implement a relighting interface using GaussianShader [JTL23] to achieve higher resolution and incorporate transparent fluid rendering within the original 3DGS CUDA rasterizer.
4.1 Velocity Optimization and Reconstruction
Figure 10 compares four types of $V_{vol}$: (1) velocity estimated from the input video and used as optimization guidance, (2) a simulation with random physical parameters, and two optimized simulations employing either (3) partial or (4) complete loss terms as defined in Eq. 9. This is also used as an ablation study to verify both the necessity and efficacy of our optimization approach. Specifically, we examine the extent to which the optimization changes the parameters and how these changes influence the simulation outcomes, particularly in terms of the resulting 3D velocity fields. Besides, as shown in Fig. 10, the reconstructed dynamics can even correct artifacts present in the video-estimated velocity fields through physical simulation.
Figure 10: Visualization of (a) volumetric velocity obtained from the video, (b) simulation result of randomly initialized physical parameters, (c) simulation result of optimized parameters with only MSE loss, and (d) simulation result of optimized parameters with full loss terms.
Table 1 compares the scale variation of several parameters and the the simulation results before and after the optimization process. Starting from arbitrary initializations at plausible scales, the optimized parameters consistently converge to stable values. This convergence is expected, as these values represent the optimum for the given reconstructed terrain and estimated 3D velocity. In contrast, manual estimation of parameter scales often fails to produce simulation results that match the velocity field extracted from the video. As shown in Fig. 10 (b), simulations using the same geometry but without proper optimization exhibit significant fluid overflow beyond the terrain. Furthermore, even initialized with scales comparable to the final optimized values, the subtle discrepancies in the magnitude can lead to pronounced differences in the simulation over time, ultimately making the dynamic reconstruction fail.
[…]
Table 1: Optimization effects on typical physical parameters and simulated results for the reconstruction scale as the case in Fig. 10.
Two synthetic videos generated from known fluid simulation data are used to compare the reconstructed 2D velocity fields against the ground truth. Since our method inputs planar velocity (discretized by frame rate and scaled to screen dimensions), both the ground truth and reconstructed 3D velocities are projected into 2D NDC space for visual comparison. Figure 11 presents the result. The reconstructed in-plane velocity captures mesoscale dynamics but retains noise on fluid surfaces. By optimizing physics-based simulation and aligning velocity directions, we achieve smoother results while enhancing physical details—such as boundary collisions in the river flow and strain development in the honey’s inflow region.
Figure 11: Comparison of planar velocity for fluid as river (top) and honey (bottom). Column (a) is the ground truth, column (b) is the estimated planar velocity in early stage of the framework as described Fig.1, and column (c) is the result by our method.
4.2 User Study
We conducted a user study to perform a qualitative evaluation, comparing the results of our method with those generated by TriplaneGaussian across all frames of the input videos (videos of the results are included in the supplementary material). The study included the 8 examples in Fig. 3D Dynamic Fluid Assets from Single-View Videos with Generative Gaussian Splatting and Fig. 9. Users were asked to evaluate the outputs along three dimensions: (1) similarity to the input video, (2) realism of the output, and (3) aesthetic quality. For each dimension, participants rated the results on a scale from 1 to 5, with higher scores indicating better performance. We invited 15 participants, all of whom are either full-time 3D artists or graduate students specializing in 3D design. The scores from the user study are visualized in Fig. 12, where our method achieves noticeably higher scores overall.
Figure 12: Plot of the scores from the user study. Left: scores for TriplaneGaussian. Right: scores for our results. Higher scores, indicated by the greener regions, represent better performance.
4.3 Evaluation on Different Fluid Types
We apply our method to two different types of liquid fluids – inviscid and viscous liquid – to show the generalization of our framework in Fig. 9. The effects of gravity and boundary conditions are more pronounced in inviscid fluids with well-defined shapes (Fig. 9 (a) to (e)). For viscous fluids (Fig. 9 (f) to (g)), the relative motion between different parts is constrained by viscosity and further dampened by the boundary. By optimizing additional simulation parameters (e.g., viscosity), the resimulated fluid can mimic a wide range of dynamics observed in the video. Meanwhile, the rocks on the water surface (Fig.3D Dynamic Fluid Assets from Single-View Videos with Generative Gaussian Splatting, Fig.9 (a) and (e)) can also be successfully detected, benefiting from the pre-processing of the generated 3DGS.
4.4 Performance with Different 3DGS Generation Methods
As a generic framework based on generative 3DGS, our method is not dependent on a specific method of Gaussian generation. We test our method with different 3DGS generation methods: generative methods, TriplaneGaussian (Fig. 9 (a), (c) to (g)) and TRELLIS [XLX24] (Fig. 9 (b)). These methods differ in terms of image fidelity, resolution, and creativity, thus catering to varying user needs in real-world applications. TriplaneGaussian benefits from triplane encoding of image textures, which aligns well with the input image, though this comes at the cost of lower resolution. TRELLIS shows more divergence from the input images, even at high Classifier Free Guidance (CFG) scales, but produces Gaussians at a higher resolution compared to TriplaneGaussian.
4.5 User Editing
@@
For users looking to edit the generated fluid asset, the framework provides easy modification of simulation parameters to achieve different fluid effects, change the fluid’s material, adjust the downstream flow, or introduce solid-liquid interactions as shown in Fig. 3D Dynamic Fluid Assets from Single-View Videos with Generative Gaussian Splatting and Fig. 13. Our output asset also includes a generated terrain, which makes the fluid easily concatenated to and produce interaction with existing 3D scenery. In this way, fluid behavior within scenic videos can be faithfully reproduced in the digital world.
Figure 13: Examples of our output after user editing. (a): Several interacting objects are added, extending the downstream length of the fluid. (b): The appearance of the fluid is modified, and the gravity direction is changed. (c): Fluid texture can be flexible, generated terrain is twisted, and reconstructed dynamics is interactive.
4.6 Runtime
@@
Figure 14: Runtime (in seconds, on a logarithmic scale) of different components of the algorithm for the example shown in Fig. 3D Dynamic Fluid Assets from Single-View Videos with Generative Gaussian Splatting.
5 Conclusions
In this paper, we introduced a novel two-stage pipeline for reconstructing consistent and accurate fluid dynamics from a single-view video, starting with geometry generation and motion reconstruction and then optimizing simulation parameters. Beyond liquid fluids, our method can potentially be extended to gaseous fluids, with necessary enhancements to be explored in future work. Our experiments show that output results of viscous fluids can be directly rendered using the reconstructed 3DGS model. In contrast, for inviscid fluids with more amorphous features, rendering by relighting yields surface textures more consistent with the flow motion. This approach has opened up new possibilities in the field of fluid asset creation. Found by the experiment, our approach currently needs certain adaptions for real use. The quality of the output dynamic 3D fluids would depend on the grid resolution used for optimization and simulation, for input video with flows not that much small-scale dynamics, our appraoch produces satisfcaoty reconsrtuction with afforadble computational cost. but if those videos has complex and detailed splashing, our appraoch needs quite high-res grid-scale and smaller time-step to optimze and simualte, put real-applciation not plausible.
Our experimental results indicate that the practical application of our approach might require certain adaptations. The quality of the reconstructed dynamic 3D fluid dependents on the grid resolution used for optimization and simulation. For input videos dominated by large-scale flow with minimal small-scale dynamics, our method produces satisfactory reconstructions at an affordable computational cost. However, for videos featuring complex, detailed phenomena such as splashing, a higher grid resolution and a smaller simulation time-step are necessary. These requirements can make the computational cost prohibitively high for real-time or interactive applications. Meanwhile, it currently cannot handle long videos with high-torrent fluids due to increased complexity and convergence issues in simulation parameter optimization. Color changes not related to velocity (e.g., in lava due to temperature) may also affect the accuracy of our optical-flow-based velocity estimation before extra parameters are added to the optimization. Future work will aim to address these limitations for more accurate reconstructions.